OpenStudy (anonymous):

PLEASE HELP ME !!!!!!!!!!!!!!!!!!!!!!!!11 Turner can read 4/5 of a chapter in 11 minutes. How many chapters can he read in 37 minutes

jimthompson5910 (jim_thompson5910):

4/5 = 8/10 (multiply top and bottom by 2) so 4/5 = 0.8

jimthompson5910 (jim_thompson5910):

This means he can read 0.8 of a chapter in 11 minutes

jimthompson5910 (jim_thompson5910):

Let x = # of chapters he can read in 37 minutes if we set up a proportion, we get 0.8/11 = x/37 Cross multiply to get 0.8*37 = 11x 29.6 = 11x Now divide both sides by 11 to isolate x 29.6/11 = x 2.690909090909 = x x = 2.690909090909 ... this is the decimal form of the answer x = 148/55 ... convert to a fraction if needed